If the pair of lines ax2 + 2(a + b)xy + by2 = 0 lie along diameters of a circle and divide the circle into four sectors such that the area of one of the sectors is thrice the area of another sector then ?

Grade:12

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
From the given data
\frac{2\sqrt{(a+b)^{2}-ab}}{a+b}=1
(a+b)2=4(a2+b2+ab)
3a2+3b2+2ab=0
Thanks and Regards
Shaik Aasif
askIITians faculty